(67/14) – (24/14) =                      .
(67/14) – (24/14) =                      .
(67/14) – (24/14) =                      .

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student


(67/14)–(24/14)
= 43/14.
Fractions with same denominators are called like fractions.
Difference of two fractions = (67 - 24)/14

= 43/14


Thanks